OpenStudy (anonymous):

Breann solved the system of equations below. What mistake did she make in her work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 20 x = 4 2(4) + y = 5 y = −3

OpenStudy (anonymous):

Well, If i'm Not Mistaken Breann Did Not Make A Mistake...She Did Everything Right.

OpenStudy (anonymous):

let me see.... 2x + y = 5 x - 2y = 10 y = 5 - 2x (she used the 1st equation and solved for y) correct x - 2(5 - 2x) = 10 (she subbed 5 - 2x in for y in 2nd equation) correct x - 10 + 4x = 10 (she multiplied through the parenthesis) correct 5x - 10 = 10 (she combined like terms) correct 5x = 20 (she got 5x by itself) correct x = 4 (she solved for x) correct 2(4) + y = 5 (she subbed 4 in for y in the 1st equation) correct 8 + y = 5 (she skipped this step but you wouldn't need to write this) y = - 3 (she solved for y) correct check...(I am just doing this for myself) x - 2y = 10 4 - 2(-3) = 10 4 + 6 = 10 10 = 10 (correct) There is nothing wrong with Breann's answer. It is correct. By the way....wow....what a username u got (lol)....what made you decide to pick that one ....ROFL